# CVBench: Benchmarking Cross-Video Synergies for Complex Multimodal Reasoning

### v3

@@ 5 Conclusion
CRediT authorship contribution statement
Nannan Zhu: Writing, Visualization, Validation, Software, Methodology, Conceptualization; Yonghao Dong: Software, Methodology; Teng Wang: Writing, Investigation; Xueqian Li: Data curation, Software; Shengjun Deng: Investigation, Data curation, Software; Yijia Wang: Investigation, Data curation; Zheng Hong: Investigation, Data curation; Tiantian Geng: Data curation, Investigation; Guo Niu: Validation, Supervision, Resources, Project administration; Hanyan Huang: Investigation, Funding acquisition, Formal analysis; Xiongfei Yao: Investigation, Data curation; Shuaiwei Jiao: Investigation, Data curation.
Declaration of competing interest
The authors declare no conflict of interest.
Data availability
The benchmark dataset and evaluation code are available at: https://github.com/Hokhim2/CVBench.
Acknowledgment
This work is supported by the Guangdong Basic and Applied Basic Research Foundation (2022A1515140103), the Research Projects of Ordinary Universities in Guangdong Province (2023KTSCX133), and the Guangdong Ordinary University Key Laboratory (2021KSYS008).
Appendix A Task definitions
@@
While single-video tasks assess a model’s comprehension of individual videos, CVBench evaluates cross-video relational reasoning. This critical capability enables MLLMs to advance from controlled lab settings to robust real-world deployment. CVBench introduces 15 carefully curated evaluation dimensions that collectively measure a model’s capacity for inter-video relational reasoning, each targeting a unique but complementary aspect of cross-video understanding. An overview of these tasks, their concepts, differences, and example questions is provided in Table LABEL:long_label.
@@
Task 1: Cross-video anomaly detection
Concept: The comparative analysis of multi-source video data identifies anomalous segments deviating from normal patterns through global spatiotemporal feature and behavioral pattern comparisons.
Difference: Analyzes anomaly saliency through multi-video event richness and viewpoint coherence rather than single-video perspectives.
Example questions:
Question: Which video features individuals directly engaged in construction work rather than showcasing materials or interior design? Option: A. Video 1, B. Video 2, C. Video 3, D. Video 4
Task 2: Cross-video scene recognition Concept: Fusion of semantic features from multiple videos to locate scene segments matching target spatiotemporal attributes. Difference: Focuses on cross-video differences rather than single-video similarities. Example questions:
Question: Did both videos feature a team successfully scoring a decisive goal or point during the match’s closing moments? Option: Yes/No.
Task 3: Multi-video key-action recognition
Concept: This category involves spatiotemporal alignment of motion trajectories across videos to identify differences in action execution, supporting action quality assessment and optimization.
Difference: It emphasizes modeling the spatiotemporal consistency of actions across multiple videos, rather than classifying actions within a single video.
Example questions:
Question: What is the key action difference between the presenter in Video 1 and the presenter in Video 2? Option: A. In Video 1, holding products is key, while in Video 2, gesturing with hands dominates, B. In Video 1, gesturing is emphasized more, while in Video 2, interacting with products is dominant, C. In both, sitting is a primary action, but outdoor walking is unique to Video 2, D. Video 2 uses textual prompts for interactions, while Video 1 does not.
Task 4: Cross-video event retrieval
Concept: This category involves rapidly locating segments across multiple videos that meet specific event elements, enabling effective filtering and selection.
Difference: The events are distributed across different videos, rather than being contained within a single video.
Example questions:
Question: Which video shows the celebration of a team after successfully leading or winning in the match?
Option: A. Video 1, B. Video 2, C. Both of them, D. None of them.
Task 5: Cross-video object recognition
Concept: This category involves fusing multi-view object features to address challenges such as occlusion and deformation, enabling consistent identity recognition of objects across videos.
Difference: It leverages multi-source information to compensate for the limitations of single-video perspectives.
Example questions:
Question: Which item is commonly recognized across both Videos 1 and 2 despite differences in design aesthetic?
Option: A. Bamboo plants, B. Distressed wooden door, C. Umbrella stands, D. Metalwork tools.
Task 6: Multi-video attribute recognition
Concept: This category involves confirming and extracting the attributes (such as texture, color, function, relationship, characteristic and shape) of a specific target across multiple videos, capturing the same target in different states.
Difference: It focuses on cross-video attribute state transitions, rather than static attribute recognition within a single frame.
Example questions:
Question: What color predominantly features in the attire of winning players across both sporting events?
Option: A. White, B. Red, C. Blue, D. Green.
Task 7: Joint-video counting Concept: This category involves the precise identification and statistical analysis of the same target across multiple videos. Difference: The targets are distributed across multiple videos rather than a single video, requiring cross-video unified authentication of the targets, while eliminating limitations of single-video perspectives (such as occlusion). Targets may also be abstract, such as events. Example questions:
Question: How many monitors other than cell phones are in all the videos? Option: A. 1, B. 2, C. 3, D. 4.
Task 8: Cross-video entity matching Concept: This category involves making similarity judgments of entities across multiple videos with varying spatiotemporal conditions (different space and time, same space and different times, different space and different times), such as identifying criminal suspects.
Difference: It emphasizes identity association across non-overlapping fields of view, overcoming the limitations of single-view tracking. Example questions:
Question: Which brand is consistently featured in both videos, reinforcing the high-performance gaming theme? Option: A. ASUS, B. Corsair, C. Gigabyte, D. RGB.
Task 9: Multi-view scene understanding
Concept: This category involves integrating spatiotemporal clues from multiple perspectives to reconstruct the complete causal chain and logical relationships of event development.
Difference: It constructs a global event knowledge graph, rather than understanding event fragments from a single perspective.
Example questions:
Question: What is the common theme across the videos that ties together the movie date, romantic dinner, family daily, and cobblestone alley walk? Option: A. Solitude, B. Companionship, C. Adventure, D. Conflict.
Task 10: Multi-video temporal reasoning
Concept: This category involves integrating multiple videos and making judgments about hidden logical relationships at specific times, such as predicting the future or sorting events.
Difference: It addresses the asynchronous timestamp issue in single videos, enabling global temporal modeling.
Example questions:
Question: Based on the interconnections, what is the correct chronological order of the criminal case depicted across the videos? Option: A. 1-2-3-4, B. 4-1-2-3, C. 1-2-4-3, D. 3-1-4-2.
Task 11: Joint-video spatial navigating
Concept: This category involves fusing multi-view geometric information to construct a 3D spatial semantic map, supporting cross-view path planning.
Difference: Spatial registration and joint reasoning of multi-source visual data.
Example questions:
Question: Where are the girl’s two wardrobes located in the room across all videos? Option: A. Head to the top-right corner to enter the kitchen and stove is located in the top-right corner of the kitchen, B. Head to the top-left corner to enter the kitchen and stove is located in the top-left corner of the kitchen, C. Head to the
Task 12: Video difference caption
Concept: This category involves fine-grained cross-video comparison, identifying differences across multiple videos in dimensions such as event progression and object states.
Difference: It emphasizes the identification of differences in dynamic processes, rather than static image comparison.
Example questions:
Question: What differentiates Video 1’s organization method from Video 2 and Video 3? Option: A. Uses a rotating tray, B. Combines different gadgets, C. Focuses on vertical space, D. Emphasizes utensil storage.
Task 13: Cross-video counterfactual reasoning
Concept: Based on the spatiotemporal factual foundation from multiple videos, this category constructs a causal inference chain for a virtual scenario.
Difference: Causal effect estimation based on multi-source observational data, overcoming the limitations of single-video causal inference.
Example questions:
Question: What element remains unchanged across the videos, regardless of the setting? Option: A. Clothing style, B. Dance type, C. Footwear focus, D. Presence of music.
Task 14: Joint-video summarization
Concept: This category involves extracting semantic information from multiple videos to generate event logic and descriptions that satisfy specific query requirements.
Difference: Cross-video information fusion and semantic compression, rather than simple fragment stitching.
Example questions:
Question: How can the overall narrative of Videos 1, 2, and 3 be summarized?
Option: A. A guide on photo gallery exploration, cloud synchronization, and DSLR photography tips, B. A sequential explanation of transferring photos, securing them, and organizing them within a smartphone, C. Three independent tutorials on iPhone photo management, Google Photos app features, and manual gallery browsing, D. A step-by-step guide on the physical, cloud, and app-based photo storage and transfer across platforms.
Task 15: Cross-video procedural transfer Concept: This category focuses on cross-video dynamic knowledge transfer capabilities, this approach analyzes the knowledge in the source video (such as first aid steps or mathematical problem solving methods) and adaptively refines or supplements the execution logic of similar tasks in the target video, while overcoming the interference caused by scene differences. Difference: Associating procedural semantics across videos, as a single video can only capture isolated processes and cannot achieve knowledge generalization or adaptation. Example questions:
Question: What decor themes from Video 2 can be transferred to Video 4 to enhance the bar’s elegance? Option: A. Floral arches and greenery, B. Heart motifs on walls, C. Outdoor nature-inspired pieces, D. Grand tents and fairy lights.
Appendix B QA generation
We propose a hierarchical framework for generating cross-video question-answer pairs with three key stages for comprehensive evaluation. In the first stage, adaptive video segmentation is combined with GPT-4o to generate fine-grained spatiotemporal annotations, capturing critical elements such as object interactions and action sequences. The prompt template for generating cohesive video summaries is shown in Table 8. In the second stage, a relationship extraction module constructs a cross-video association graph, identifying multidimensional relationships such as object evolution and event continuity. The cross-video relation extraction prompt is depicted in Table 9. Based on 15 professional templates, the system generates diverse questions, ranging from basic matching to advanced reasoning, ensuring quality through multi-model validation and difficulty calibration. The QA pair prompt template is shown in Table 10.
Table 8: Prompt template for generating cohesive video summaries with GPT-4o.
You are an expert in analyzing videos. You will receive the video consisting of a series of frames. Your task is to carefully observe the details of the video frames and generate a detailed and accurate description of a video in English based on your observations. Instructions for writing the detailed description:
a). Focus on the key visual details such as appearance, motion, scene, actions, objects involved, and interactions between elements in the video.
b). Pay attention to the type of video and the important points like the order of events, appearance and actions of people or objects, and any significant changes, events or movements.
c). Note every detail in the video and give a clear and easy to understand description.
d). Indicators like ([Video number], Timestamp: [start time]-[end time]/[total time]), for example, (Video 1, Timestamp: 00:00-00:10/01:20), mean that the video is divided into parts, and the timestamps represent the start time and end time of the part. Please note that the change in Video number means that a new video has been inputed, the description of the new video has to be restarted, and the flow between the different parts of the same video has to ensure narrative continuity and avoid abrupt breaks in the description.
e). There is no need for markdown titles or emphasis such as ”**XXX**” or words such as ”(Frame X)” or ”(Video X, Timestamp: 00:XX-00:XX/0X:XX)”, emphasize the time in a uniform way, using a simple (00:XX-00:$X$) will do!
Table 9: Cross-video relation extraction prompt. Analyze the interconnections between multiple videos by combining their captions.
You are an expert in analyzing interconnections of multiple videos. We extracted detailed captions (pertimestamp) for each video. For example, ’Video 1 : caption: 00:00-00:10: corresponds to the caption generated for the first video between 00:00 and 00:10.
Your task is to generate a detailed and accurate summary of interconnections between the videos (Video 1, Video2, etc.) as a paragraph, based on all the given video captions and videos. Make sure not to lose any important information. Use the following details to create a clear and complete narrative: Instructions for writing the detailed summary:
a). Pay strong attention to the interconnections between the videos, including differences, correlations, matches, etc. For instance, the correlations can be that the videos are split by a complete video, just in a different order; The differences can be different types of videos, or the same objects in them doing different actions; Matches can be the same person appearing in different videos. Check if events across videos follow a logical sequence, identify shared locations or objects across videos and note complementary angles.
b). Identify the theme of each video and use your intelligence to learn to infer hidden links.
c). Note the commonalities and contradictions between videos. The timestamps are just more detailed to allow you to understand the complete video, the focus is between the videos and not between the timestamps.
d). Do not mention that the information comes from captions.
e). Also, use your common sense to judge interconnections between the videos.
Table 10: QA pair prompt. Create quiz questions based on the interconnections between video captions.
I would like you to act as a quizmaster who designs questions based on provided captions of each videos and summary of the interconnections between videos that should be grounded in real-life contexts and made more relevant to everyday realities. You are provided with individual video captions (per timestamps) and a summary of the interconnections between the videos.
You are to generate relevant QA pairs by analyzing key interconnections between videos while also referring to the detailed video summary. Your questions should specifically target the interconnections between the videos. Your question should be answerable, and checkable.Guidelines for Response:
1. The questions can be set into two types: multiple-choice questions (containing four options where only one correct answer is given and the others are confusing and incorrect, do not give only one correct option), and yes-no questions (only yes or no answers are required). If descriptions are ambiguous, (e.g., timeline or content), prioritize answer uniqueness. Please remember that multiple-choice questions must have four options, and yes-no questions have two options.
2. Avoid excessive similarity between options, and be more careful to avoid answers that can be inferred directly from the options.
3. When involving multiple subjects, use subject attributes (e.g.,“adult in red”) for reference. Avoid abstract identifiers like “subject 1” or “person A”.
4. Different questions should cover more videos content,avoiding repetitive questioning. 4. Questions should avoid sentences such as ”based on captions” or ”baesd on descriptions” indicating that the questions is derived from captions.
5. Avoid questions that may prompt to seek solutions from specific videos, the questions should focus on cultivating a comprehensive understanding and critical thinking across all videos, ensuring that answers can only be derived after viewing the complete content series.
6. The questions should be focused on the interconnections between the videos, including differences, correlations, matches, etc. Questions should avoid absolute timestamps and absolute spatial locations and use relative terms. For example, after drinking water, next to the TV, etc.
7. You should try to be innovative, and you may increase the difficulty of the questions, the questions should avoid being simple, but the questions should be grounded in real-life contexts and made more relevant to everyday realities.
This framework integrates visual content analysis with logical reasoning, featuring a multi-layered quality control system that includes multi-model consensus validation and hallucination detection to ensure the verifiability and distinctiveness of the generated questions. Ultimately, the framework provides a comprehensive evaluation spectrum, ranging from single-hop to commonsense reasoning, with multidimensional metrics such as accuracy and reasoning depth. Table 8, 9, and 10 present standardized prompt templates, ensuring the systematization and reproducibility of the evaluation process.
Appendix C More CVBench examples
CVBench’s 15 carefully designed subcategories of videos and questions evaluate key dimensions such as spatiotemporal correlation, entity-event association, and causal-logical reasoning. These tasks cover various challenges, from tracking object consistency across cameras to inferring causal relationships between multiple videos, detecting spatial annotation conflicts, and performing cross-modal data fusion. Each subcategory presents complex tasks that closely resemble real-world scenarios. More examples of these tasks are provided in Fig. 6.

### v4

@@ 5 Conclusion
Appendix A Dataset Collection
A.1 Task Definitions
@@
While single-video tasks assess a model’s comprehension of individual videos, CVBench evaluates cross-video relational reasoning. This critical capability enables MLLMs to advance from controlled lab settings to robust real-world deployment. CVBench introduces 15 carefully curated evaluation dimensions that collectively measure a model’s capacity for inter-video relational reasoning, each targeting a unique but complementary aspect of cross-video understanding. An overview of these tasks, their concepts, differences, and example questions is provided in Table 7.
Task 1: Cross-video anomaly detection Concept: The comparative analysis of multi-source video data identifies anomalous segments deviating from normal patterns through global spatiotemporal feature and behavioral pattern comparisons. Difference: Analyzes anomaly saliency through multi-video event richness and viewpoint coherence rather than single-video perspectives. Example Questions: Question: Which video features individuals directly engaged in construction work rather than showcasing materials or interior design? Option: A. Video 1, B. Video 2, C. Video 3, D. Video 4 Task 2: Cross-video scene recognition Concept: Fusion of semantic features from multiple videos to locate scene segments matching target spatiotemporal attributes. Difference: Focuses on cross-video differences rather than single-video similarities. Example Questions: Question: “Did both videos feature a team successfully scoring a decisive goal or point during the match’s closing moments?” Option: Yes/No. Task 3: Multi-video key-action recognition Concept: This category involves spatiotemporal alignment of motion trajectories across videos to identify differences in action execution, supporting action quality assessment and optimization. Difference: It emphasizes modeling the spatiotemporal consistency of actions across multiple videos, rather than classifying actions within a single video. Example Questions: Question: What is the key action difference between the presenter in Video 1 and the presenter in Video 2? Option: A. In Video 1, holding products is key, while in Video 2, gesturing with hands dominates., B. In Video 1, gesturing is emphasized more, while in Video 2, interacting with products is dominant., C. In both, sitting is a primary action, but outdoor walking is unique to Video 2., D. Video 2 uses textual prompts for interactions, while Video 1 does not.
Task 4: Cross-video event retrieval Concept: This category involves rapidly locating segments across multiple videos that meet specific event elements, enabling effective filtering and selection. Difference: The events are distributed across different videos, rather than being contained within a single video. Example Questions: Question: Which video shows the celebration of a team after successfully leading or winning in the match?
Option: A. Video 1.", "B. Video 2.", "C. Both of them.", "D. None of them.
Task 5: Cross-video object recognition Concept: This category involves fusing multi-view object features to address challenges such as occlusion and deformation, enabling consistent identity recognition of objects across videos. Difference: It leverages multi-source information to compensate for the limitations of single-video perspectives. Example Questions: Question: Which item is commonly recognized across both Videos 1 and 2 despite differences in design aesthetic? Option: A. Bamboo plants.", "B. Distressed wooden door.", "C. Umbrella stands.", "D. Metalwork tools. Task 6: Multi-video attribute recognition Concept: This category involves confirming and extracting the attributes (such as texture, color, function, relationship, characteristic and shape) of a specific target across multiple videos, capturing the same target in different states. Difference: It focuses on cross-video attribute state transitions, rather than static attribute recognition within a single frame. Example Questions: Question: What color predominantly features in the attire of winning players across both sporting events? Option: "A. White.", "B. Red.", "C. Blue.", "D. Green." Task 7: Joint-video counting Concept: This category involves the precise identification and statistical analysis of the same target across multiple videos. Difference: The targets are distributed across multiple videos rather than a single video, requiring cross-video unified authentication of the targets, while eliminating limitations of single-video perspectives (such as occlusion). Targets may also be abstract, such as events. Example Questions: Question: How many monitors other than cell phones are in all the videos? Option: A. 1, B. 2, C. 3, D. 4. Task 8: Cross-video entity matching Concept: This category involves making similarity judgments of entities across multiple videos with varying spatiotemporal conditions (different space and time, same space and different times, different space and different times), such as identifying criminal suspects.
Task 10: Multi-video temporal reasoning Concept: This category involves integrating multiple videos and making judgments about hidden logical relationships at specific times, such as predicting the future or sorting events. Difference: It addresses the asynchronous timestamp issue in single videos, enabling global temporal modeling. Example Questions: Question: Based on the interconnections, what is the correct chronological order of the criminal case depicted across the videos? Option: A. 1-2-3-4.", "B. 4-1-2-3.", "C. 1-2-4-3.", "D. 3-1-4-2. Task 11: Joint-video spatial navigating Concept: This category involves fusing multi-view geometric information to construct a 3D spatial semantic map, supporting cross-view path planning. Difference: Spatial registration and joint reasoning of multi-source visual data. Example Questions: Question: Where are the girl’s two wardrobes located in the room across all videos? Option: A. Head to the top-right corner to enter the kitchen and stove is located in the top-right corner of the kitchen, B. Head to the top-left corner to enter the kitchen and stove is located in the top-left corner of the kitchen, C. Head to the Task 12: Video difference caption Concept: This category involves fine-grained cross-video comparison, identifying differences across multiple videos in dimensions such as event progression and object states. Difference: It emphasizes the identification of differences in dynamic processes, rather than static image comparison. Example Questions: Question: What differentiates Video 1’s organization method from Video 2 and Video 3? Option: A. Uses a rotating tray., B. Combines different gadgets., C. Focuses on vertical space., D. Emphasizes utensil storage. Task 13: Cross-video counterfactual reasoning Concept: Based on the spatiotemporal factual foundation from multiple videos, this category constructs a causal inference chain for a virtual scenario. Difference: Causal effect estimation based on multi-source observational data, overcoming the limitations of single-video causal inference. Example Questions: Question: What element remains unchanged across the videos, regardless of the setting? Option: A. Clothing style., B. Dance type., C. Footwear focus., D. Presence of music. Task 14: Cross-video object recognition Concept: This category involves fusing multi-view object features to address challenges such as occlusion and deformation, enabling consistent identity recognition of objects across videos. Difference: It leverages multi-source information to compensate for the limitations of single-video perspectives. Example Questions: Question: Which item is commonly recognized across both Videos 1 and 2 despite differences in design aesthetic? Option: A. Bamboo plants.", "B. Distressed wooden door.", "C. Umbrella stands.", "D. Metalwork tools. Task 15: Cross-video event retrieval Concept: This category involves rapidly locating segments across multiple videos that meet specific event elements, enabling effective filtering and selection. Difference: The events are distributed across different videos, rather than being contained within a single video. Example Questions: Question: Which video shows the celebration of a team after successfully leading or winning in the match?
Option: A. Video 1.", "B. Video 2.", "C. Both of them.", "D. None of them.
@@
A.2 QA Generation
We propose a hierarchical framework for generating cross-video question-answer pairs with three key stages for comprehensive evaluation. In the first stage, adaptive video segmentation is combined with GPT-4o to generate fine-grained spatiotemporal annotations, capturing critical elements such as object interactions and action sequences. The prompt template for generating cohesive video summaries is shown in Figure 6. In the second stage, a relationship extraction module constructs a cross-video association graph, identifying multidimensional relationships such as object evolution and event continuity. The cross-video relation extraction prompt is depicted in Figure 7. Based on 15 professional templates, the system generates diverse questions, ranging from basic matching to advanced reasoning, ensuring quality through multi-model validation and difficulty calibration. The QA pair prompt template is shown in Figure 8.
Figure 6: Prompt template for generating cohesive video summaries with GPT-4o. generate a cohesive summary of the following video by merging the segment-wise captions into a comprehensive temporal description that fully encapsulates the video’s content.
This framework integrates visual content analysis with logical reasoning, featuring a multi-layered quality control system that includes multi-model consensus validation and hallucination detection to ensure the verifiability and distinctiveness of the generated questions. Ultimately, the framework provides a comprehensive evaluation spectrum, ranging from single-hop to commonsense reasoning, with multidimensional metrics such as accuracy and reasoning depth. Figures 6, 7, and 8 present standardized prompt templates, ensuring the systematization and reproducibility of the evaluation process.
Figure 7: Cross-video relation extraction prompt. Analyze the interconnections between multiple videos by combining their captions.
Figure 8: QA pair prompt. Create quiz questions based on the interconnections between video captions.
A.3 More CVBench Examples Visualization
CVBench’s 15 carefully designed subcategories of videos and questions evaluate key dimensions such as spatiotemporal correlation, entity-event association, and causal-logical reasoning. These tasks cover various challenges, from tracking object consistency across cameras to inferring causal relationships between multiple videos, detecting spatial annotation conflicts, and performing cross-modal data fusion. Each subcategory presents complex tasks that closely resemble real-world scenarios. More examples of these tasks are provided in Figure 9.